If y varies inversely as x and y = 725 when x = 20, find x when y is 50.
Answer by Alan3354(69443)

If y varies inversely as x and y = 725 when x = 20, find x when y is 50.
------------
y = k/x
725 = k/20
k = 14500
==============
50 = 14500/x
50x = 14500
x = 290